# Learning Neural Control Barrier Functions from Offline Data with Conservatism

### v1

@@ 5.2 Results
Table 1: Evaluation results for different pairs of nominal controllers and safety filters. Results of each pair are averaged across 500 runs for the 2D navigation task and 50 runs for the vision-based navigation task. Green: Safest agent with the highest average success % across all controllers.
@@
Success % Collision % Success % Collision % Success % Collision % Success % Collision %
@@
CCBF-based filters improve safety significantly while minimally affecting rewards
@@
Observing the results in Table 1 for the 2D navigation task, we can see that NCBF and CCBF result in similar success percentages and both result in zero collisions, outperforming iDBF. In the vision-based navigation environment, CCBF outperforms both NCBF and iDBF by achieving better success and collision percentages using any nominal controller. Unlike the case of the 2D navigation scenario, iDBF generally performs better than NCBF. The improved performance of iDBF can be attributed to the naturally increased distance between training trajectories, which reduces the likelihood of mislabeling nearby images as safe or unsafe as the sampled images nearby a training trajectory will not overlap with the images visited by other training trajectories. In contrast, the training trajectories in the 2D environment are more closely spaced, resulting in the sampled states by iDBF, which are labeled as unsafe, overlapping with the states of training trajectories which are labeled as safe.
@@
In the Swimmer environment, CCBF reduces the average cost over all nominal controllers by approximately 60% with around a 6% drop in average reward. In contrast, both iDBF and NCBF increase the cost. In the Hopper environment, both CCBF and iDBF exhibit similar behavior seen in the Swimmer environment, whereas NCBF reduces the average cost by 47% with only a 7% drop in average reward, which differs from its performance in the Swimmer environment. These results show that CCBF consistently reduces cost while minimally impacting reward. On the other hand, NCBF’s results vary between environments—performing poorly in Swimmer but well in Hopper—possibly because Swimmer results in a larger distribution shift during deployment.
That said, FISOR results in the lowest cost in the Swimmer environment. It also almost results in the lowest cost in the Hopper environment—comparable to BC-safe and only slightly less safe than COptiDICE. However, it underperforms in terms of reward achieved in both environments. In the Swimmer environment, it results in a reward of -0.08, whereas CCBF paired with BC-Safe results in a 0.45 reward. Similarly, in the Hopper environment, its reward return is 0.08 compared to 0.17 resulting from CCBF paired with the COptiDICE-based nominal controller.
CCBF is easier to train and more robust to hyperparameters
@@
In all experiments, CCBFs consistently distinguished safe and unsafe states when using different values for $w_{c}$ in the loss term $\mathcal{L}_{c}$. Figure 4(a) (in Appendix D) visualizes the CCBFs learned using different values of $w_{c}$ for the 2D navigation example. All of them classify the states in the obstacle as unsafe and the states visited by the safe trajectories as safe. However, tuning $w_{c}$ and $\tau$ is needed for CCBF to better classify OOD states as unsafe. On the other hand, iDBF’s performance seems to be more sensitive to the density threshold $p$ that is used to classify sampled actions as OOD. It is often the case that a choice of $p$, especially when it is large (e.g., $p$ = 0.1), requires the tuning of the other hyperparameters $w_{\text{safe}}$, $w_{\text{unsafe}}$, $\varepsilon_{\text{safe}}$, and $\varepsilon_{\text{unsafe}}$ to prevent training an iDBF that assigns a single value to all states. As can be seen in Figure 4(b) (in Appendix D), without the right choice of $p$, iDBF does not properly classify the safe, unsafe, InD, and OOD states. We observed similar behavior of iDBF in the vision-based navigation and safety gymnasium scenarios.
6 Conclusions and Limitations
We introduced Conservative Control Barrier Functions (CCBFs), neural safety filters trained using offline data. The idea is to add a loss term that penalizes over-optimistic safety estimations of OOD states to the loss function for training neural CBFs, resulting in safety estimations that better differentiate safe, unsafe, and InD and OOD states. When used to define quadratic programs as safety filters, they result in closed-loop behaviors that avoid unsafe and OOD states. We evaluated our approach in multiple offline environments and showed that it outperforms existing baselines in maintaining safety while minimally affecting performance.
Our approach in its present form does not offer formal guarantees that it will maintain safety or OOD states avoidance, and deriving such guarantees would be an important future direction. Moreover, evaluating generated safety filters in more realistic environments including physical robots operating in real environments would be equally valuable. Also, using world models [55, 56] to model the dynamics over the latent space would be needed to model more complex environments that are partially-observed. Finally, using dynamic programming-like approaches for training the safety filters might be needed when the states in the offline datasets are not labeled as safe and unsafe, but functions that define failure sets are available [21, 57].
Appendix A Additional details on experimental setup
A.1 Environments
(a) 2D Navigation
(b) Vision-based Nav
(c) SafetyHopper
(d) SafetySwimmer
Figure 2: Illustration of the four experimental environments used in our evaluation.
A.2 Experimental details for 2D navigation environment
A.2.1 Dataset generation and implementation details
We generated 2,000 diverse trajectories using a CBF-QP policy that combines a nominal PD controller with a manually-designed CBF. The nominal controller is defined as $\mathbf{u}_{\text{nom}}=K_{p}(\mathbf{x}_{g}-\mathbf{x})+K_{d}\frac{\mathbf{u}_{\text{nom}}}{\Delta t}$, where $\mathbf{x}_{g}$ is the goal position, $\Delta t=0.1$ seconds is the timestep (10 Hz frequency), and $K_{p},K_{d}$ are the proportional and derivative gains, respectively. The constraints in the QP are $\dot{h}(\mathbf{x})+\alpha h(\mathbf{x})\geq 0$, where $h(\mathbf{x})=\|\mathbf{x}-\mathbf{x}_{\text{obs}}\|^{2}-r^{2}$ is the CBF, and $-v_{\text{max}}\leq v_{x}\leq v_{\text{max}}$ and $-v_{\text{max}}\leq v_{y}\leq v_{\text{max}}$, where $v_{\text{max}}=3$. Each trajectory is initialized from a random state. We gradually increase obstacle radius $r$ as the dataset generation progresses. Initially, we set r to $0.01$ and increment it by 0.2 up to 5, updating every 25000 time steps, where the time step count is cumulative over trajectories. The episode ends after 200 timesteps or when the agent reaches the goal, defined as $\|\mathbf{x}-\mathbf{x}_{\text{g}}\|^{2}<\delta$, where $\delta$ is 0.5. This setup generates a diverse dataset of safe and unsafe trajectories. We train the CCBF as described in Section 4.3, but we do not include the terms $\mathcal{L}_{\text{descent}}$ and $\mathcal{L}_{\text{lip}}$ in this scenario.
During evaluation, each episode starts from a randomly sampled initial state within the square $[-18,5]^{2}$. We solve the QP without actuation constraints, and then scale the control input as $\mathbf{u}_{\mathrm{scaled}}=\frac{\mathbf{u}}{\|\mathbf{u}\|_{2}}\,v_{\max}$ to ensure it satisfies the actuation constraints.
A.2.2 Visualization of a trajectory generated using CCBF
Figure 3: A trajectory generated using CCBF‑QP in the 2D navigation task as well as the nominal and CCBF‑QP controls.
A.2.3 Hyperparameters
Table 3: Hyperparameters for NCBF, CCBF, iDBF in the 2D navigation task
Parameter Value
Number of hidden layers 3
Hidden dimension 128
Batch size 128
Learning rate $1\times 10^{-4}$
$\varepsilon_{\text{safe}}$ 0.2
$\varepsilon_{\text{unsafe}}$ 0.2
$w_{\text{safe}}$ 1.0
$w_{\text{unsafe}}$ 1.2
$w_{\text{ascent}}$ 1.0
$p$ $1\times 10^{-8}$
$w_{c}$ 0.5
$\tau$ 0.7
Optimizer Adam
Time step (dt) 0.1
A.3 Experimental details for vision-based navigation environment
A.3.1 Dataset generation and implementation details
Using the same setup as in the 2D navigation task, we set $\delta=2$ and collect 3,000 trajectories with $64\times 64\times 3$ RGB image observations. They are a mix of safe and unsafe trajectories, generated by varying the obstacle radius and initial positions as before, with the maximum episode length being 20 seconds (200 timesteps, 10Hz controller). The collected image-based trajectories are then used to learn latent dynamics, as described in Section 4.1. An observation is considered safe if the Euclidean distance $d$ to the obstacle is greater than or equal to $r+\epsilon$, where $\epsilon=3$, otherwise it is unsafe. We train the CCBF as described in Section 4.3, but do not include $\mathcal{L}_{\text{lip}}$.
A.3.2 Hyperparameters
Table 4: Hyperparameters for NCBF, CCBF, and iDBF in the vision-based navigation task
Parameter Value
Latent dimension size 4
Action dimension size 2
Hidden dimension 128
Number of hidden layers 3
Batch size 256
Learning rate $1\times 10^{-4}$
$\varepsilon_{\text{safe}}$ 0.08
$\varepsilon_{\text{unsafe}}$ 0.15
$\varepsilon_{\text{ascent}}$ 0.02
$\varepsilon_{\text{descent}}$ 0.02
$w_{\text{ascent}}$ 2.0
$w_{\text{decent}}$ 1.0
$w_{\text{safe}}$ 1.0
$w_{\text{unsafe}}$ 1.1
$p$ $1\times 10^{-6}$
$w_{c}$ 1
$\tau$ 0.7
Optimizer Adam
Time step (dt) 0.1
Table 5: Hyperparameters for latent dynamics in the vision-based navigation task
Parameters Value
Encoder CNN layers
Conv2d(3$\rightarrow$32)
Conv2d(32$\rightarrow$64)
Conv2d(64$\rightarrow$128)
Flatten
Linear(8192$\rightarrow$400)
Linear(400$\rightarrow$4) (latent state)
Decoder CNN layers
Linear(4$\rightarrow$8192)
Unflatten (128, 8, 8)
ConvTranspose2d(128$\rightarrow$64)
ConvTranspose2d(64$\rightarrow$32)
ConvTranspose2d(32$\rightarrow$3)
Sigmoid activation (output $\in[0,1]$)
Input channels 3
Latent dimension size 4
Hidden dimension (Dynamics ) 400
Number of hidden layers (Dynamics) 3
Learning rate (Dynamics) $1\times 10^{-4}$
Learning rate (encoder, decoder) $1\times 10^{-4}$
Batch size 32
$\lambda_{1}$ 1.0
$\lambda_{2}$ 1.0
$\lambda_{3}$ 0.5
Optimizer Adam
Time step (dt) 0.1
A.4 Safety gymnasium Swimmer and Hopper environments
A.4.1 Implementation and experimental details
We first train a dynamics model as described in Section 4.1, but instead of employing an autoencoder as in the vision-based task to map observations to latent states, we learn the dynamics directly in safety gymnasium’s native state space. Then, we train the CCBF as described in Section 4.3, but do not include $\mathcal{L}_{\text{descent}}$. We train the nominal controllers using the offline RL algorithms (COptiDICE [28], BEAR-L [58, 27], BCQ-L [58, 59]) and using BC variants as implemented in DSRL [51]. During evaluation, if the QP was infeasible, we select the closest point in $\mathcal{U}$ to the constraint boundary.
A.4.2 Normalized reward and normalized cost metrics in the safety gymnasium tasks
For both the Swimmer and the Hopper scenarios, they can can be modeled using Constrained Markov Decision Processes (CMDP). We assume that the transition functions in these CMDPs are deterministic and in the form of the control-affine dynamics in equation (4). Each episode results in a trajectory $\sigma:=\{x_{1},u_{1},r_{1},c_{1},\dots,x_{T},u_{T},r_{T},c_{T}\}$, where the subscript represents the index of the sampling time.
The normalized reward and cost returns of a trajectory $\sigma$ generated using policy $\pi$ are defined as follows: $R_{\text{normalized}}^{\pi}:=\frac{R_{\sigma}^{\pi}-R_{\text{min}}}{R_{\text{max}}-R_{\text{min}}}\times 100$ and $C_{\text{normalized}}^{\pi}:=\frac{C_{\sigma}^{\pi}+\epsilon}{\kappa+\epsilon}$, where $R_{\sigma}^{\pi}=\sum_{t}r_{t}$ denotes the policy’s reward return when generating trajectory $\sigma$, $C_{\sigma}^{\pi}=\sum_{t}c_{t}$ denotes its cost return, and $R_{\text{min}}$ and $R_{\text{max}}$ denote the maximum and minimum empirical reward returns over all trajectories in the dataset as defined and computed in DSRL [51]. In all benchmark tasks, the cost $c_{t}$ is zero if the state $x_{t}$ is safe and one if it is unsafe.
The parameter $\kappa$ represents the cost threshold used to train the safe offline RL model, while $\epsilon$ is a small positive constant introduced to keep $C_{\text{normalized}}^{\pi}$ well-defined when $\kappa=0$. We set $\kappa=20$ and $\varepsilon=0$ in all of our experiments. In the case of FISOR, $\kappa$ is not used for training the policy and only used to compute the normalized cost. We say that a trajectory is safe when the normalized cost, $C_{\text{normalized}}$, satisfies the condition $C_{\text{normalized}}<1$, implying that the cost return is less than the specified threshold when $\varepsilon=0$.
A.4.3 Hyperparameters
Table 6: Hyperparameters for the NCBF and dynamics model in the safety gymnasium environments
Parameter Value
hidden dimension (CBF) 128
Number of hidden layers (CBF) 4
Learning rate (CBF) $1\times 10^{-5}$
$\varepsilon_{\text{safe}}$ 0.08
$\varepsilon_{\text{unsafe}}$ 0.15
$\varepsilon_{\text{ascent}}$ 0.02
$w_{\text{ascent}}$ 2
$w_{\text{safe}}$ 1.0
$w_{\text{unsafe}}$ 1.2
$w_{\text{lip}}$ 1.0
Batch size 256
Hidden dimension (dynamics) 128
Number of hidden layers (dynamics) 4
Learning rate (dynamics) $1\times 10^{-4}$
Optimizer Adam
Time step (dt) 0.1
For each of Swimmer and Hopper, we selected the best hyperparameters across three different models when reporting averaged results. In the Hopper environment, we choose the best $w_{c}$ and $\tau$ parameters for the CCBF models as $(0.1,1)$, $(0.05,1)$, and $(0.5,0.7)$, respectively. For the iDBF models in Hopper, we select $p$ values of $1\times 10^{-10}$, $1\times 10^{-5}$, and $1\times 10^{-8}$, respectively. Similarly, in the Swimmer environment, we select the best $w_{c}$ and $\tau$ parameters for the three CCBF models as $(1,0.5)$, $(1,1)$, and $(1,0.7)$, respectively. For the iDBF models in Swimmer, we select $p$ values of $1\times 10^{-6}$, $1\times 10^{-9}$, and $1\times 10^{-12}$, respectively. We also train the safety filters in Swimmer and Hopper for 15k and 50k steps respectively and choose the best checkpoints.
Appendix B Ablation studies
In this section, we present two ablation experiments designed to analyze: (1) the performance of iDBF for different model sizes and density thresholds $p$, and (2) the effect of labeling all OOD states as unsafe when training a NCBF from offline data, as opposed to our method which simply lowers the safety score of these OOD states relative to the InD state from which they were sampled around, without the need to label them as unsafe.
B.1 Influence of model size and density threshold $p$ on performance of iDBF
To remain faithful to the iDBF methodology and address whether model size may be causing the sub-par closed-loop performance observed in the safety gymnasium environments, we evaluated two iDBF architectures. The first is a medium-sized neural network with 4 hidden layers of 128 neurons each, and the second is a larger model with 5 hidden layers of 400 neurons each. Tables 7 and 8 report the corresponding results.
@@ B.1 Influence of model size and density threshold $p$ on performance of iDBF
B.2 Learning a NCBF from data labeled using an extremely conservative strategy
This ablation examines whether explicitly labeling OOD states as unsafe (as in iDBF) yields better results than softly lowering their safety scores relative to the InD safe state from which they were sampled around (as in CCBF). We used an extreme labeling approach: from each safe state in the dataset, we sampled random actions, took a one‑step forward simulation, and marked all the resulting states as unsafe. This procedure mimics the effect of choosing a very high density threshold $p$ within the iDBF framework. We call this approach Extreme iDBF. Results are shown in Table 9.

### v2

@@ V-B Results
TABLE I: Controller evaluation results (S=Success%, C=Collision%). Results averaged across 500 runs (2D) and 50 runs (vision-based). Green: Safest agent with the highest average success % across all controllers.
@@
S C S C S C S C
@@
CCBF-based filters improve safety while minimally affecting reward returns
Observing the results in Table I for the 2D navigation task, we can see that NCBF and CCBF result in similar success percentages and both result in zero collisions, outperforming iDBF. Trajectories taken when using CCBF and NCBF as a safety filter are presented in Figure 1.
@@
In the vision-based navigation environment, CCBF outperforms both NCBF and iDBF by achieving better success and collision percentages using any nominal controller. Unlike the case of the 2D navigation scenario, iDBF generally performs better than NCBF. The improved performance of iDBF can be attributed to the naturally increased distance between training trajectories, which reduces the likelihood of mislabeling nearby images as safe or unsafe as the sampled images nearby a training trajectory will not overlap with the images visited by other training trajectories. In contrast, the training trajectories in the 2D environment are more closely spaced, resulting in the sampled states by iDBF, which are labeled as unsafe, overlapping with the states of training trajectories which are labeled as safe.
@@
In the Swimmer environment, CCBF reduces the average cost over all nominal controllers by approximately 60% with around a 6% drop in average reward. In contrast, both iDBF and NCBF increase the cost. In the Hopper environment, both CCBF and iDBF exhibit similar behavior seen in the Swimmer environment, whereas NCBF reduces the average cost by 47% with only a 7% drop in average reward, which differs from its performance in the Swimmer environment. These results show that CCBF consistently reduces cost while minimally impacting reward. On the other hand, NCBF’s results vary between environments, performing poorly in Swimmer but well in Hopper, possibly because Swimmer results in a larger distribution shift during deployment.
That said, FISOR results in low costs in the Swimmer and Hopper tasks (comparable to CCBF paired with BC-safe in the Swimmer task and with COptiDICE in the Hopper task). However, FISOR yields much lower rewards in both environments: -0.08 in Swimmer compared to 0.45 in the case of CCBF paired with BC-Safe, and 0.08 in Hopper compared to 0.17 when CCBF is paired with COptiDICE.
It is worth noting that the policies used during evaluation in Table II are distinct from those used to collect the DSRL dataset. This demonstrates that CCBFs, despite being trained on offline data, can effectively improve safety for new policies without noticeably affecting task performance.
iDBF performs well with the BC-Safe controller, improving safety in vision-based navigation and maintaining comparable levels in Swimmer and Hopper. This is expected since iDBF’s approach labels states as unsafe when they correspond to low-probability actions under BC-Safe. However, with other controllers such as COptiDICE, performance degrades, likely because iDBF steers them toward regions that BC-Safe deems safe but that are suboptimal if the dataset trajectories are not expert (reward-maximizing).
To further investigate iDBF’s performance, we conducted extensive ablation experiments by varying the model size and the density threshold $p$ for selecting OOD actions. We tested medium-sized models (4 layers, 128 neurons) versus large models (5 layers, 400 neurons) across different density thresholds ($p\in\{1,0.1,10^{-4},10^{-8}\}$). For medium-sized models, average costs increased substantially across all thresholds compared to not using any safety filter: average Swimmer costs rose from 1.49 to 4.38-6.92, while Hopper costs increased from 0.74 to 1.02-2.50. Large models generally showed improvement with average costs ranging from 3.85-6.63 for Swimmer and 0.77-1.90 for Hopper, demonstrating better performance than medium-sized models across most density thresholds. However, across all configurations and model sizes, iDBF’s performance consistently deteriorated and underperformed compared to nominal controllers, and to CCBF and NCBF methods that leverage actual unsafe trajectories in the offline dataset. Detailed results of these experiments can be found in Appendix A-A and A-B.
CCBFs is easier to train and more robust to changes in hyperparameters
@@
In all experiments, CCBFs consistently distinguished safe and unsafe states when assigning $w_{c}$ to values between 0 and 0.5. However, tuning $w_{c}$ (typically between 0.5 and 1) was needed for CCBF to sufficiently disincentivize unseen actions, as shown in Figure 1. On the other hand, iDBF’s performance was more sensitive to the density threshold $p$ that is used to classify sampled actions as OOD. We also had to tune the hyperparameters $w_{\text{safe}}$, $w_{\text{unsafe}}$, $\varepsilon_{\text{safe}}$, and $\varepsilon_{\text{unsafe}}$ to obtain an iDBF that does not assign a single value to all states.
Isolating the effect of our conservativeness-inducing loss term $\mathcal{L}_{c}$
We conducted an experiment where we trained CCBFs with $w_{c}\in\{0,0.5,1.0\}$, for the vision-based navigation, Swimmer, and Hopper environments. Note that when $w_{c}=0$, our approach is equivalent to training a NCBF. We split the dataset trajectories into training and testing ones and only train on the former. For each safe state in the test trajectories, we evaluated the CCBF at the next state along the trajectory as well as at a set of states obtained by taking random actions from that same state, following the procedure used to compute $\mathcal{L}_{c}$. First, we found that increasing $w_{c}$ decreases the CCBF values of the states reached by taking both the actions in the test set and the actions randomly sampled, as expected. Second, increasing $w_{c}$ enlarges the gap between the values of the CCBFs at the states in the test set compared to the ones reached by the random actions. This clearly shows that the learned CCBFs generalize well to distinguish actions taken by the policies generating the data and random ones and consider the states resulting from the latter as less safe. In contrast, we found that NCBFs consistently assign similar values for the randomly-reached states and the ones in the test set, also as expected. We present examples of our results in Figure 2 and show more detailed ones in Figures 3, 4, and 5.
Next, we studied whether other conservativeness-inducing loss terms would result in similar improvement in performance compared to the NCBF approach. We denote the first candidate loss term, which replaces $\mathcal{L}_{c}$, by $\mathcal{L}_{\text{unsafe}}^{\prime}$ and its associate weight by $w_{\text{unsafe}}^{\prime}$. The term $\mathcal{L}_{\text{unsafe}}^{\prime}$ is equivalent $\mathcal{L}_{\text{unsafe}}$ evaluated at the states sampled in the same manner when computing $\mathcal{L}_{c}$, i.e., treating the states reached by randomly-sampled actions similarly to unsafe ones. We call this new approach CCBF. It exhibits similar properties to CCBF when comparing the values it assigns for the OOD and InD states in new trajectories and performs better than iDBF and NCBF, but under performs compared to CCBF. We summarise the main results below and refer readers to Appendix A-C for details. For Swimmer, the average results across all controllers are: CCBF with $w_{c}^{*}=0.5$ ($C=1.33$, $R=0.34$) and $w_{c}^{*}=1$ ($C=2.44$, $R=0.32$). For Hopper, the averages are: CCBF with $w_{c}^{*}=0.5$ ($C=0.37$, $R=0.31$), and $w_{c}^{*}=1$ ($C=0.91$, $R=0.26$). We note that the learned CCBF* resulting from $w_{c}^{*}=1$ under performs the one resulting from $w_{c}^{*}=0.5$. This potentially occurs because $w_{c}^{*}=1$ penalizes OOD states as much as actual unsafe states. Since randomly reached states are likely (but not necessarily) OOD, this aligns with findings in [43] stating that treating OOD and unsafe actions similarly degrades performance. As a remark, in both CCBF and CCBF*, increasing $w_{c}$ and $w_{c}^{*}$ above the value of one can result in over-conservative safety filters which lead to low rewards.
From these two experiments, we conclude that CCBF’s and CCBF*’s conservativeness-inducing loss terms enable learning safety filters that not only filter actions for safety, but also for staying in-distribution. Our experiments show that CCBF outperforms CCBF*, but further evaluation is needed to claim that one of these approaches is consistently better than the other, which we leave for future work.
Figure 2: Variation of the mean of the barrier values of states reached by taking either dataset actions (orange) or randomly sampled actions (green) starting from safe states in unseen test trajectories, plotted against the number of training steps. Results are shown for CCBF trained with $w_{c}=0.5$: (a) Swimmer, (c) Hopper, and (e) vision-based navigation, where dataset actions are consistently evaluated as safer than random actions. For NCBF: (b) Swimmer, (d) Hopper, and (f) vision-based navigation, where this distinction is not observed and the two curves largely overlap.
VI Conclusion
We introduced Conservative Control Barrier Functions (CCBFs), neural safety filters trained using offline data. The core idea is to add a loss term that prevents over-estimation of the safety of state-action pairs not seen in the dataset. When reference controllers and safety filters are trained offline, CCBFs, even when paired with reference controllers that were not used to generate the offline dataset, outperform baselines and result in closed-loop behaviors that better preserve safety while minimally impacting task performance. Our approach in its present form does not offer formal guarantees for safety maintenance or OOD state avoidance, and deriving ones would be exciting future work. Also, evaluating our method on maintaining the safety of real robots would also be needed for further validation.
Appendix A Ablation studies
A-A Influence of model size and density threshold $p$ on performance of iDBF
To remain faithful to the iDBF methodology and address whether model size may be causing the sub-par closed-loop performance observed in the safety gymnasium environments, we evaluated two iDBF architectures. The first is a medium-sized neural network with 4 hidden layers of 128 neurons each, and the second is a larger model with 5 hidden layers of 400 neurons each. Tables VIII and IX report the corresponding results.
@@ A-A Influence of model size and density threshold $p$ on performance of iDBF
A-B Learning iDBF with density threshold $p=1$
From each state in the dataset comprised of safe trajectories, we sampled random actions, took a one‑step forward simulation, and marked all the resulting states as unsafe. This procedure mimics the effect of choosing a very high density threshold $p=1$ within the iDBF framework. We call this approach Extreme iDBF. Results are shown in Table X.